OpenStudy (anonymous):

the average of 1/5 and other two numbers out of which one is the half of other is 1/4 the smallest fraction out of unknown is?

OpenStudy (akshay_budhkar):

you asked it again?

OpenStudy (anonymous):

yes i don't understand

OpenStudy (akshay_budhkar):

ok

OpenStudy (akshay_budhkar):

do you know what is the average of 3 numbers?

OpenStudy (anonymous):

converting to decimal (.2 + x + x/2)/3 = .25 .2 + x + x/2 = .75 x(1+(1/2)) = .55 x(1.5) = .55 converting back to fractions x(3/2) = (11/20) x = (2/3) * (11/20) x = 11/30 x/2 = 11/60 which is the smaller of the fractions

OpenStudy (anonymous):

If you need me to explain any of the steps please ask